# Carve3D: Improving Multi-view Reconstruction Consistency for Diffusion Models with RL Finetuning

### 21 Dec 2023 v1

@@ 3.1 Evaluate Consistency via NeRF Reconstruction
NeRF [31] is widely adopted as the 3D representation for learning text-to-3D tasks. A 3D model represented by NeRF can be reconstructed from the view images of the object and their corresponding camera poses. The quality of a NeRF notably depends on the consistency of the provided images images [31, 52] – inconsistent views lead to artifacts in the NeRF, which includes floaters, blurring, and broken geometry. To address this challenge, we introduce a metric for assessing the consistency among multiple views.
The intuition behind MRC comes from the relationship between multi-view consistency and the reconstructed NeRF. As shown in Figure 3, when the multi-view images are consistent, they can produce a well reconstructed NeRF, preserving almost all the visual cues from the input images; therefore, the views rendered from the NeRF at the same camera viewpoints will look the same as the original views; conversely, when the multi-view images are inconsistent (e.g., intentionally introduced inconsistency in Figure 3), they will produce a NeRF with broken geometry and floater artifacts; thus, the NeRF rendered views will look different from the original views. Building upon this observation, we propose the MRC metric, defined as the image distances between the original multi-view images and the views of the reconstructed NeRF rendered at the same viewpoints, as illustrated in Figure 2.
@@ 3.2 Implementation
We formulate the implementation of MRC as three parts: fast sparse-view NeRF reconstruction, measuring image distance between the input images and the rendered images, and a normalization technique for the image distance. The pseudo code for our MRC implementation is shown in References in Appendix.
@@
We conduct NeRF reconstruction with sparse-view Large Reconstruction Model (LRM) proposed in  [24, 17]. Different from dense view NeRF reconstruction [31, 8, 33], sparse-view LRM reconstructs a NeRF with only $4$-$6$ view images. Also, with its feed-forward reconstruction, it can achieve a speed two orders of magnitude faster than previous optimization-based reconstruction methods. MRC leverages all multi-view images for both NeRF reconstruction and 3D consistency evaluation. Although the NeRF is reconstructed based on the visual prior of the input multi-views images, the rendering from the same views still exhibits notable differences if there is inconsistency inside the input, as shown in Figure 3.
@@
In Section 3.1, the consistency problem is reduced from 3D to a 2D image dissimilarity problem. To measure the image dissimilarity between the input views and their corresponding NeRF rendered views, we utilize the perceptual image distance metric, LPIPS [56]. LPIPS exhibits smoother value changes with respect to the consistency of multi-view images compared to PSNR, SSIM, L1, and L2, as shown in Figure 14 in Appendix). Such smoothness is derived from the non-pixel-aligned computation in LPIPS, as opposed to the other image distance metrics that are more pixel-aligned. Also, the smoothness is a crucial aspect for MRC to serve as the reward function in RLFT, because non-smooth, high-variance reward functions makes the RLFT training more challenging.
@@
Current multi-view diffusion models [46, 24, 28, 58] target single object generation with background. Consequently, if computing LPIPS on the entire image, trivially reducing the object’s relative size (as illustrated in Appendix  Figure 9’s car example) can exploit MRC, as the majority of images will be the white background. Therefore, we propose normalizing our metric with respect to the object’s size. Specifically, we identify the smallest square bounding box of the foreground object in the input view image. Then we crop both the input images and the rendered images with that bounding box, resize them to a fixed resolution, and evaluate the LPIPS. This normalization effectively prevents the reward hacking of MRC by diminishing foreground object sizes, as shown in Figure 9 in Appendix.
@@ 3.3 Metric Experiment
Figure 3 shows the qualitative result on one object of our MRC metric experiment. With increased inpainting area size, the NeRF rendered view also shows larger image difference, which is then captured by MRC’s image distance metric, LPIPS. Figure 4 presents the quantitative curve of the same experiment. MRC indeed shows a monotonically increasing pattern as the views become more inconsistent. As shown in Figure 14, MRC constantly exhibits monotonically increasing pattern, and it is also smoother than the other MRC variants using PSNR, SSIM, L1, and L2. For metric experiments on other distortion types, see Appendix D.
@@ 4 RLFT for Multi-view Consistency
In the previous section, we proposed a fast and reliable multi-view consistency metric, and in this section we describe how it can be used to finetune a multi-view diffusion model. We propose RLFT for enhancing the consistency of 2D multi-view diffusion models, using the negative MRC as the reward function (Figure 2). Building upon DDPO [5], we opt for its pure on-policy policy gradient algorithm over the default partially on-policy version for substantially improved training stability. To maintain proximity to the base model, we incorporate KL divergence regularization similar to  [15, 36]. In addition, we scale up the RLFT to achieve higher rewards by studying the scaling laws [20] of diffusion model RLFT through extensive experiments.
@@ 4.1 Preliminaries on DDPO
$s_{t}$ $=(c,t,x_{t}),$
$a_{t}$ $=x_{t-1},$
$\pi(a_{t}|s_{t})$ $=p_{\theta}(x_{t-1}|c,t,x_{t}),$
$r(s_{t},a_{t})$ =$=\begin{cases}r(x_{0},c)&\text{if }t=0,\\ 0&\text{otherwise},\end{cases}$
$r(x_{0},c)$ $=-\text{MRC}(x_{0})$
where each denoising step is a timestep, $c$ is the context, i.e. the text prompt, $x_{t}$ is the image being denoised at step $t$, $p_{\theta}$ is the diffusion model being finetuned, $x_{T}$ is the initial noisy image, $x_{0}$ is the fully denoised image, and $r(x_{0},c)$ is the negative MRC (References in Appendix) computed on the fully denoised image.
@@
In order to optimize the model with respect to the reward function, a family of RL algorithms, known as policy gradient methods, are commonly adopted, such as REINFORCE [53] and Proximal Policy Optimization (PPO) [45]. $\text{DDPO}_{\text{SF}}$ is based on the vanilla policy gradient algorithm, REINFORCE [53], also known as the Score Function (SF). On the other hand, $\text{DDPO}_{\text{IS}}$ builds upon PPO [45] and conducts multiple optimization steps per round of data using an importance sampling (IS) estimator and importance weight clipping.
As a common practice to reduce the variance of the policy gradients [32], DDPO [5] uses the advantages, which are rewards normalized to have zero mean and unit variance, instead of directly using the rewards. Specifically, the mean and standard deviation statistics of the rewards are tracked for each prompt $c$:
@@
DDPO’s normalizing advantage replaces the value model that is more widely adopted in PPO-based RLHF methods [36, 54, 50]. This is similar to the recent work [25], which shows that the value model creates unnecessary computation cost that can be replaced with a simpler advantage formulation.
By using the advantage term (Equation 1) in place of the reward, the $\text{DDPO}_{\text{SF}}$ policy gradient function is:
@@
where the expectation is taken over data generated by the policy $\pi_{\theta}$ with the parameters $\theta$. The log probability $\log p_{\theta}(x_{t-1}|c,t,x_{t})$ can be easily obtained since the policy is an isotropic Gaussian distribution when using the DDIM sampler [5, 47]. The $\text{DDPO}_{\text{IS}}$ (Equation 6 in Appendix) function has an additional importance sampling term than Equation 2.
Black et al. [5] choose $\text{DDPO}_{\text{IS}}$ as the default policy gradient function, because it exhibits better sample efficiency than $\text{DDPO}_{\text{SF}}$ (Fig. 4 of [5]). Such choice is consistent with the use of PPO [45] in LLM RLHF literature [36, 54, 50, 2, 3].
Figure 5: Reward curves on the testing set with 4 different random seeds for IS and SF versions, where negative MRC is used as the reward. The IS version produces reward curves with high variance, including two runs that fail to improve the reward and collapse. In contrast, the SF version stably produces reward curves with low variance.
Figure 6: When we only plot KL divergence without incorporating KL regularization, we observe qualitative correlation between the KL value and the prompt alignment degradation. Despite being distant in the finetuning process, epoch 15 and epoch 45, which have lower KL divergence to the base model, generates prompts better aligned with the prompts. On the other hand, epoch 30, which has much higher KL divergence to the base model, generates results with broken identity, i.e. the body of the pug is missing.
@@ 4.2 Improvements over DDPO
While RLFT using the default $\text{DDPO}_{\text{IS}}$ coupled with MRC can enhance the 3D consistency of multi-view diffusion models, it still faces challenges regarding training stability, the shift of output distributions, and an unclear training scale setting to achieve optimal rewards with minimal distribution shift. To address these issues, we propose three improvements over DDPO [5] in this section. Given the universal nature of these challenges in RLFT, our enhancements may offer broader applicability across various tasks.
@@ 4.2.1 Pure On-policy Training
Training stability is a major challenge in both RLFT [7, 59] and traditional RL [14]. With the default $\text{DDPO}_{\text{IS}}$, our training process is evidently unstable, as shown in Figure 5. Training experiments with different random seeds or a slight change of hyperparameters can lead to different reward curves and qualitative results. This complicates the training result evaluation as we cannot distinguish meaningful improvement or deterioration from the variance introduced by random seed.
We argue that such high variance is derived from the multi-step update in $\text{DDPO}_{\text{IS}}$ [5], originally proposed in PPO [45]. While it theoretically allows for better sample efficiency similar to off-policy methods [45], it also causes the training to be more unstable and the reward curves to be more variant, because it uses data collected with the older policy to update the newer policy. Due to the undesirable consequences of training instability, we adopt the pure on-policy variant $\text{DDPO}_{\text{SF}}$, discarding the multi-step update from PPO (Equation 6 in Appendix). As shown in Figure 5, $\text{DDPO}_{\text{SF}}$ significantly improves the training stability of our RLFT, while maintaining a comparable sample efficiency as the default $\text{DDPO}_{\text{IS}}$.
Diverging from DDPO [5] and most LLM RLHF literature [36, 54, 50, 2, 3], we choose REINFORCE [53] ($\text{DDPO}_{\text{SF}}$) over PPO [45] ($\text{DDPO}_{\text{IS}}$) for its superior training stability. We provide two hypotheses behind our surprising finding in Section B.3, including the difficulty of the task reward function and the size of the model being finetuned. The favored use of REINFORCE [53] over PPO [45] could apply to broader scenarios that meet these two conditions. We leave the verification of our hypotheses as future work.
@@ 4.2.2 KL Divergence Regularization
In RLFT methods, distribution shift (also known as reward overoptimization) can lead to low-quality results, such as cartoon-like, less realistic style [5] or oversaturated colors and unnatural shape [15], despite achieving high rewards. In our case, we observe this as degradation of diversity, texture details and prompt alignment after prolonged RLFT with the MRC reward. Previous methods [36, 15] suggest mitigating reward overoptimization by incorporating a penalty on the KL divergence between the log probabilities of the outputs from the base and the finetuned models. In our case, the base model is Instant3D-10K [24] without any additional finetuning. By plotting the KL divergence values during finetuning, we also find that KL divergence correlates to the reward overoptimization problem (Figure 6), suggesting us to adopt KL divergence regularization.
@@
Our advantage terms now consist of both the normalized reward and the normalized KL divergence. Our final policy gradient function, used in our experiments, is a combination of Equations 2 and 4
@@
Figure 7: Scaling law for Carve3D’s diffusion model RLFT algorithm. When we scale up the amount of compute for RLFT, the model improves its reward smoothly under the optimal data size. The amount of compute scales linearly with respect to the batch size. The reward curves also become more stable (less variant) with a larger batch size. The reward curves are reported up to epoch $50$.
@@ 4.2.3 Scaling Laws for RLFT
The training of RL is highly sensitive to the chosen scale setting [14], impacting various results, including the final converged reward. Through the scaling laws identified from extensive experiments, we scale up the amount of compute (equivalent to scaling up the batch size in our case) to achieve the optimal reward. Although our scaling experiments are only conducted with the multi-view consistency task, our insights into the scaling laws of diffusion model RLFT are general and can be beneficial in broader scenarios.
@@
The model’s reward does not directly scale with the data size but there exists a more complicated relationship between them. As shown in Figure 7, the optimal data size at each batch size grows as the batch size get larger, indicating that both factors need to be scaled up in tandem; after the optimal data size, naively continuing to scale up the data size actually reduces the model’s reward. Surprisingly, even when trained on a prompt set as small as a size of $10$, the model still shows generalization to the testing prompts. We choose data size of 30 with batch size 768 in our final experiments as it achieves the highest reward in our analysis.
@@
With the pure on-policy $\text{DDPO}_{\text{SF}}$ (REINFORCE [53]), the model steadily and smoothly improves its rewards throughout the finetuning process, meaning that more training epochs constantly lead to higher reward. However, from our qualitative results, we also observe worse distribution shift, e.g. the degradation of prompt alignment and texture details, as training epoch increases. Due to the correlation between KL divergence and the quality degradation (Figure 6), we stop the finetuning early when a predefined KL divergence threshold is reached. This threshold is empirically chosen based on qualitative results. For fair comparisons, we report the reward curves up to epoch 50 in Figure 7. See Section B.1 for the definition of epoch in RLFT, which is different from its definition in supervised learning.

### 09 Apr 2024 v2

@@ 3.1 Evaluate Consistency via NeRF Reconstruction
A 3D model represented by Neural Radiance Field (NeRF) can be reconstructed from the view images of the object and their corresponding camera poses. The quality of a NeRF notably depends on the consistency of the provided images images [31, 53] – inconsistent views lead to artifacts in the NeRF, which includes floaters, blurring, and broken geometry. To address this challenge, we introduce a metric for assessing the consistency among multiple views.
The intuition behind MRC comes from the relationship between multi-view consistency and the reconstructed NeRF. As shown in Fig. 3, when the multi-view images are consistent, they can produce a well reconstructed NeRF, preserving almost all the visual cues from the input images; therefore, the views rendered from the NeRF at the same camera viewpoints will look the same as the original views; conversely, when the multi-view images are inconsistent (e.g., intentionally introduced inconsistency in Fig. 3), they will produce a NeRF with broken geometry and floater artifacts; thus, the NeRF rendered views will look different from the original views. Building upon this observation, we propose the MRC metric, defined as the image distances between the original multi-view images and the views of the reconstructed NeRF rendered at the same viewpoints, as illustrated in Fig. 2.
@@ 3.2 Implementation
We formulate the implementation of MRC as three parts: fast sparse-view NeRF reconstruction, measuring image distance between the input images and the rendered images, and a normalization technique for the image distance. The pseudo code for our MRC implementation is shown in Appendix References.
@@
We conduct NeRF reconstruction with sparse-view Large Reconstruction Model (LRM) proposed in  [24, 17]. Different from dense view NeRF reconstruction [31, 8, 33], sparse-view LRM reconstructs a NeRF with only $4$-$6$ view images. Also, with its feed-forward reconstruction, it can achieve a speed two orders of magnitude faster than previous optimization-based reconstruction methods. MRC leverages all multi-view images for both NeRF reconstruction and 3D consistency evaluation. Although the NeRF is reconstructed based on the visual prior of the input multi-views images, the rendering from the same views still exhibits notable differences if there is inconsistency inside the input, as shown in Fig. 3.
@@
In Sec. 3.1, the consistency problem is reduced from 3D to a 2D image dissimilarity problem. To measure the image dissimilarity between the input views and their corresponding NeRF rendered views, we utilize the perceptual image distance metric, LPIPS [57]. LPIPS exhibits smoother value changes with respect to the consistency of multi-view images compared to PSNR, SSIM, L1, and L2, as shown in Appendix Fig. 14. Such smoothness is derived from the non-pixel-aligned computation in LPIPS, as opposed to the other image distance metrics that are more pixel-aligned. Also, the smoothness is a crucial aspect for MRC to serve as the reward function in RLFT, because non-smooth, high-variance reward functions makes the RLFT training more challenging.
@@
Current multi-view diffusion models [46, 24, 28, 59] target single object generation with background. Consequently, if computing LPIPS on the entire image, trivially reducing the object’s relative size (as illustrated in Appendix Fig. 9’s car example) can exploit MRC, as the majority of images will be the white background. Therefore, we propose normalizing our metric with respect to the object’s size. Specifically, we identify the smallest square bounding box of the foreground object in the input view image. Then we crop both the input images and the rendered images with that bounding box, resize them to a fixed resolution, and evaluate the LPIPS. This normalization effectively prevents the reward hacking of MRC by diminishing foreground object sizes, as shown in Appendix Fig. 9.
@@ 3.3 Metric Experiment
Fig. 3 shows the qualitative result on one object of our MRC metric experiment. With increased inpainting area size, the NeRF rendered view also shows larger image difference, which is then captured by MRC’s image distance metric, LPIPS. Fig. 4 presents the quantitative curve of the same experiment. MRC indeed shows a monotonically increasing pattern as the views become more inconsistent. As shown in Appendix Fig. 14, MRC constantly exhibits monotonically increasing pattern, and it is also smoother than the other MRC variants using PSNR, SSIM, L1, and L2. For metric experiments on other distortion types, see Appendix D.
@@ 4 RLFT for Multi-view Consistency
In Sec. 3, we proposed a fast and reliable multi-view consistency metric named MRC, and in this section we describe how it can be used to finetune a multi-view diffusion model. In this section, we present an improved Reinforcement Learning Finetuning (RLFT) algorithm for enhancing the consistency of 2D multi-view diffusion models, using the negative MRC as the reward function (Appendix References). Building upon DDPO [5], we opt for its pure on-policy version over the default partially on-policy version of the policy gradient algorithm for substantially improved training stability. To maintain proximity to the base model, we incorporate KL divergence regularization similar to  [15, 36]. In addition, we scale up the RLFT to achieve higher rewards by studying the scaling laws [20] of diffusion model RLFT through extensive experiments.
@@ 4.1 Preliminaries on DDPO
$s_{t}$ $=(c,t,x_{t}),$ (1)
$a_{t}$ $=x_{t-1},$ (2)
$\pi(a_{t}|s_{t})$ $=p_{\theta}(x_{t-1}|c,t,x_{t}),$ (3)
$r(s_{t},a_{t})$ =$=\begin{cases}r(x_{0},c)&\text{if }t=0,\\ 0&\text{otherwise},\end{cases}$ (4)
$r(x_{0},c)$ $=-\text{MRC}(x_{0})$ (5)
where each denoising step is a timestep, $c$ is the context, i.e. the text prompt, $x_{t}$ is the image being denoised at step $t$, $p_{\theta}$ is the diffusion model being finetuned, $x_{T}$ is the initial noisy image, $x_{0}$ is the fully denoised image, and $r(x_{0},c)$ is the negative MRC (Appendix References) computed on the fully denoised image.
@@
In order to optimize the model with respect to the reward function, a family of RL algorithms, known as policy gradient methods, are commonly adopted, such as REINFORCE [54] and Proximal Policy Optimization (PPO) [45]. $\text{DDPO}_{\text{SF}}$ is based on the vanilla policy gradient algorithm, REINFORCE [54], also known as the Score Function (SF) of diffusion models. On the other hand, $\text{DDPO}_{\text{IS}}$ builds upon PPO [45] and conducts multiple optimization steps per round of data using an importance sampling (IS) estimator and importance weight clipping.
As a common practice to reduce the variance of the policy gradients [32], DDPO [5] uses the advantages (Eq. 6), which are rewards normalized to have zero mean and unit variance, instead of directly using the rewards. Specifically, the mean and standard deviation statistics of the rewards are tracked for each prompt $c$:
@@
DDPO’s [5] reward-normalizing advantage replaces the value model that is more widely adopted in PPO-based [45] RLHF methods [36, 55, 51]. This is similar to [25], which shows that the value model creates unnecessary computation cost that can be replaced with a simpler advantage formulation.
By using the advantage term $A_{r}$ (Eq. 6) in place of the reward $r$, the $\text{DDPO}_{\text{SF}}$ policy gradient function is:
@@
where the expectation is taken over data generated by the policy $\pi_{\theta}$ with the parameters $\theta$. The log probability $\log p_{\theta}(x_{t-1}|c,t,x_{t})$ can be easily obtained since the policy is an isotropic Gaussian distribution when using the DDIM sampler [5, 48]. The $\text{DDPO}_{\text{IS}}$ function (Appendix Eq. 11) has an additional importance sampling term than Eq. 7.
Black et al. [5] choose $\text{DDPO}_{\text{IS}}$ as the default policy gradient function, because it exhibits better sample efficiency than $\text{DDPO}_{\text{SF}}$ (Fig. 4 of [5]). Such choice is consistent with the use of PPO [45] in Large Language Model (LLM) Reinforcement Learning from Human Feedback (RLHF) literature [36, 55, 51, 2, 3].
Figure 5: Comparing the IS and the SF versions of Carve3D reward curves on the testing set over 4 different random seeds. The IS version produces reward curves with high variance, including two runs that fails. In contrast, all runs of the SF version stably produces reward curves with low variance.
Figure 6: We observe qualitative correlation between the KL value and the prompt alignment degradation. Despite being distant in the finetuning process, epoch 15 and epoch 45, which have lower KL divergence to the base model, generates prompts that align better with the prompts. On the other hand, epoch 30, which has much higher KL divergence from the base model, generates results with broken identity, i.e. the body of the pug is missing.
@@ 4.2 Improvements over DDPO
While RLFT with the default $\text{DDPO}_{\text{IS}}$ and our MRC can enhance the 3D consistency of multi-view diffusion models, it still faces challenges regarding training stability, the shift of output distributions, and an unclear training scale setting to achieve optimal rewards with minimal distribution shift. To address these issues, we propose three improvements over DDPO [5] in this section. Given the universal nature of these challenges in RLFT, our enhancements may offer broader applicability across various tasks.
@@ 4.2.1 Pure On-policy Training
Training stability is a major challenge in both RLFT [7, 60] and traditional RL [14]. With the default $\text{DDPO}_{\text{IS}}$, our training process is evidently unstable, as shown in Fig. 5. Training experiments with different random seeds or a slight change of hyperparameters can lead to different reward curves and qualitative results. This complicates the training result evaluation as we cannot distinguish meaningful improvement or deterioration from the variance introduced by random seed.
We argue that such high variance is derived from the multi-step update in $\text{DDPO}_{\text{IS}}$ [5], originally proposed in PPO [45]. While it theoretically allows for better sample efficiency similar to off-policy methods [45], it also causes the training to be more unstable and the reward curves to be more variant, because it uses data collected with the older policy to update the newer policy. Due to the undesirable consequences of training instability, we adopt the pure on-policy variant $\text{DDPO}_{\text{SF}}$, discarding the multi-step update from PPO (Appendix Eq. 11). As shown in Fig. 5, $\text{DDPO}_{\text{SF}}$ significantly improves the training stability of our RLFT, while maintaining a comparable sample efficiency as the default $\text{DDPO}_{\text{IS}}$.
Diverging from DDPO [5] and most LLM RLHF literature [36, 55, 51, 2, 3], we choose REINFORCE [54] ($\text{DDPO}_{\text{SF}}$) over PPO [45] ($\text{DDPO}_{\text{IS}}$) for its superior training stability. We provide two hypotheses behind our surprising finding in Sec. B.3, including the difficulty of the task reward function and the size of the model being finetuned. The favored use of REINFORCE [54] over PPO [45] could apply to broader scenarios that meet these two conditions. We leave the verification of our hypotheses as future work.
@@ 4.2.2 KL Divergence Regularization
In RLFT methods, distribution shift (also known as reward overoptimization) can lead to low-quality results, such as cartoon-like, less realistic style [5] or oversaturated colors and unnatural shape [15], despite achieving high rewards. In our case, we observe this as degradation of diversity, texture details and prompt alignment after prolonged RLFT with the MRC reward. Previous methods [36, 15] suggest mitigating reward overoptimization by incorporating a penalty on the Kullback–Leibler (KL) divergence between the log probabilities of the outputs from the base and the finetuned models. In our case, the base model is Instant3D-10K [24] without any additional finetuning. By plotting the KL divergence values during finetuning, we also find that KL divergence correlates to the reward overoptimization problem (Fig. 6), suggesting us to adopt KL divergence regularization.
@@
Our advantage terms now consist of both the normalized reward (Eq. 6) and the normalized KL divergence (Eq. 9). Our final policy gradient function, used in our experiments, is a combination of Eqs. 7 and 9
@@
Figure 7: Scaling law for Carve3D’s diffusion model RLFT algorithm. When we scale up the amount of compute, the model improves its reward smoothly under the optimal data size. The amount of compute scales linearly with respect to the batch size. The reward curves also become more stable (less variant) with a larger batch size. The reward curves are reported up to epoch $50$.
@@ 4.2.3 Scaling Laws for RLFT
The training of Reinforcement Learning (RL) is highly sensitive to the chosen scale setting [14], impacting various results, including the final converged reward. Through the scaling laws identified from extensive experiments, we scale up the amount of compute (equivalent to scaling up the batch size in our case) to achieve the optimal reward. Although our scaling experiments are only conducted with the multi-view consistency task, our insights into the scaling laws of diffusion model RLFT are general and can be beneficial in broader scenarios.
@@
The model’s reward does not directly scale with the data size but there exists a more complicated relationship between them. As shown in Fig. 7, the optimal data size at each batch size grows as the batch size get larger, indicating that both factors need to be scaled up in tandem; after the optimal data size, naively continuing to scale up the data size actually reduces the model’s reward. Surprisingly, even when trained on a prompt set as small as a size of $10$, the model still shows generalization to the testing prompts. We choose data size of 30 with batch size 768 in our final experiments as it achieves the highest reward in our analysis.
@@
With the pure on-policy $\text{DDPO}_{\text{SF}}$ (REINFORCE [54]), the model steadily and smoothly improves its rewards throughout the finetuning process, meaning that more training epochs constantly lead to higher reward. However, from our qualitative results, we also observe worse distribution shift, e.g. the degradation of prompt alignment and texture details, as training epoch increases. Due to the correlation between KL divergence and the quality degradation (Fig. 6), we stop the finetuning early when a predefined KL divergence threshold is reached. This threshold is empirically chosen based on qualitative results. For fair comparisons, we report the reward curves up to epoch 50 in Fig. 7. See Sec. B.1 for the definition of epoch in RLFT, which is different from that in other contexts, and Sec. C.1 for our training data generation and curation.